OpenStudy (anonymous):

What is the missing number in this pattern? 1, 8, ___ , 64, 125

OpenStudy (anonymous):

27. 1 is 1 cubed. = 1*1*1 8 is 2 cubed = 2*2*2 27 is 3 cubed = 3*3*3 64 is 4 cubed... 125 is 5 cubed... And so on.

OpenStudy (anonymous):

thank you..

OpenStudy (anonymous):

how about this? The next two terms of the sequence 3, 6, 12 ____ are:

OpenStudy (anonymous):

That one is a bit easier. The difference between the first two integers is x or 6-3 = 3. The difference between the next two integers is y or 12-6 = 6. See the pattern? The difference between y and x is 3. So the next "difference" would be 3+3+3 = 9. Adding 9 to 12 gives us ... 21!